# On the existence of forward self-similar solutions to the two-dimensional Navier-Stokes equations

### v1

@@ 2.3. Proof of the energy estimates
since due to ${\rm div\thinspace}v_{1}=0$ and (2.21) that
${\rm div\thinspace}F(v_{0},w)={\rm div\thinspace}\left(-\Delta v_{1}-\frac{1}{2}v_{1}-\frac{1}{2}y\cdot\nabla v_{1}\right)=0.$
It follows from the Calderon–Zygmund theory that there exists a constant $c(p)$ such that
@@
$\qquad\qquad+\left.\|v_{1}\nabla v_{2}\|_{L^{p}}+\|v_{2}\nabla v_{2}\|_{L^{p}}\right)$
@@
Then from (2.26) we have
@@
The verification of this claim is a direct consequence of Lemma 2.2, and we leave it to the end of this proof. We have by using (2.35) that
@@
The application of the Gagliardo–Nirenberg interpolation inequality also gives that
@@
It remains to prove the claim of (2.35). Using (2.21), we have
$\int_{\mathbb{R}^{2}}|F(v_{0},w)|^{p}$
$\leq$ $C\int_{\mathbb{R}^{2}}\left|2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}+\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|^{p}$
$+\int_{\mathbb{R}^{2}}|\Delta w|^{p}\leq C(A,\beta,p),$
where we have used Lemma 2.2 to have for any $1<p<2$ that
$\int_{\mathbb{R}^{2}}|\Delta w|^{p}\\ \leq C(A,\beta,p),$
and that
@@
Throughout the proof, the Dirichlet energy estimate of $v_{1}$ (or equivalently Dirichlet energy estimate of $v_{0}$) (2.27) is the only place where we use the Hölder continuity exponent $\beta>0$.
We next prove Theorem 2.1 using the $H^{2}$-estimate of $v_{\text{re}}$ to be proved in Proposition 3.2.
@@
As $u_{\text{re}}(\cdot,1)=v_{\text{re}}(\cdot)$, we just prove the statement for $v_{\text{re}}$. Using the standard bootstrapping argument, along with the regularity estimates of the Stokes equations, one can show that $v_{\text{re}}\in H^{1}(\mathbb{R}^{2})$ is indeed smooth. For example, one can refer to [28, Theorem 3.8] for a detailed presentation of the bootstrapping argument. This also implies the smoothness of $u(x,1)$ as $e^{\Delta}u_{0}$ is already smooth for $t>0$. Due to the definitions of $v_{\textrm{re}}$ (2.6) and $v_{2}$ (2.19), we have
$v_{\textrm{re}}-v_{2}=v_{1}-v_{0}=(\eta_{R_{0}}-1)v_{0}+w,$
Lemmas 2.1 and 2.2 together with Theorem 2.2 tells us that for $1<p<2$,
@@
To show higher integrability, we can use estimates for elliptic equations to show that indeed $v_{\textrm{re}}$ belongs to $H^{2}(\mathbb{R}^{2})$, which will be done in Proposition 3.2. It follows from Proposition 3.2 that we have
@@
the Calderon–Zygmund estimates give that for any $p>1$ that
@@
Similarly, by Calderon–Zygmund estimates, we have
$\|\nabla q\|_{L^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
@@ 3. Weighted Energy Estimates and Pointwise Decay
In this section, we establish the pointwise decay estimates for the remainder part $v_{\text{re}}$. For the three-dimensional case, Jia and Šverák [18] first utilized localized smoothing estimates for local Leray solutions to derive pointwise decay. Here, we adopt the methods developed in [22], and establish weighted energy estimates to $\sqrt{1+|y|^{2}}v_{\text{re}}$ in $H^{2}(\mathbb{R}^{2})$, which subsequently allows us to derive the desired pointwise decay via Sobolev embeddings. That is
@@
Furthermore, we will prove the optimal decay for $v_{\text{re}}$ if the initial data $u_{0}\in C^{2}_{loc}$.
@@
(3.1) $\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0.\end{aligned}\right$.
@@
3.1. Weighted $H^{1}$-Estimates
As a first step towards the higher-order estimates, we establish a weighted $H^{1}$ bound for the remainder term $v_{\text{re}}$.
@@ 3.1. Weighted $H^{1}$-Estimates
Let the divergence-free, self-similar initial data $u_{0}(x)\in C^{0,\beta}_{loc}$ be given, and that
$A=\|u_{0}\|_{C^{0,\beta}(S^{1})}<+\infty.$
Assume $u$ is a self-similar solution to (1.1) in the sense of Definition 1.1. Consider the corresponding solution $(v_{\text{re}},q)\in H^{1}(\mathbb{R}^{2})\times L^{2}(\mathbb{R}^{2})$ to (3.1), then we have
@@
$|\nabla(\sqrt{1+|y|^{2}}v_{\text{re}})|\leq\frac{|y|}{\sqrt{1+|y|^{2}}}|v_{\text{re}}|+\sqrt{1+|y|^{2}}|\nabla v_{\text{re}}|,$
and hence
@@
As we already prove the $H^{1}$-estimates of $v_{\text{re}}$, it suffices to prove the $H^{1}$-estimates of $|y|v_{\text{re}}$.
For rigorous derivation, one should use a regularized weight $h_{\epsilon}(y)=\frac{|y|}{(1+\epsilon|y|^{2})^{1/2}}$ and test (2.20) with $\phi=|h_{\epsilon}|^{2}v_{\text{re}}$, then pass $\epsilon\to 0$. For clarity of the calculation, we present the estimate using the weight $h(y)=|y|$. We test (2.20) with $\phi=|y|^{2}v_{\text{re}}$.
@@
$=\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}-2\|v_{\text{re}}\|_{L^{2}}^{2}\geq\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}-C.$
Step 2: The Drift and Linear Damping Terms. The linear drift and damping part of the operator is $\mathcal{L}(v_{\text{re}})=-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}$. Testing this with $|y|^{2}v_{\text{re}}$, we compute:
@@
In 2D, we calculate the divergence explicitly:
@@
Combining with the damping term $\frac{1}{2}v_{\text{re}}$:
@@
This positivity is specific to the special structure of the Leray system and ensures the weighted $L^{2}$ norm is controlled.
@@
By Lemma 2.1, $|v_{0}(y)|\leq C(1+|y|)^{-1}$. Hence,
@@
Step 4: Nonlinear term $v_{\text{re}}\cdot\nabla v_{\text{re}}$ and pressure term. For the nonlinear term $v_{\text{re}}\cdot\nabla v_{\text{re}}$, we have
@@
$\int_{\mathbb{R}^{2}}\nabla q\cdot|y|^{2}v_{\text{re}}$ $=-\int_{\mathbb{R}^{2}}2qv_{\text{re}}\cdot y\leq\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}+C$
Step 5: The source terms. The source term decay fast enough to be integrable against $|y|^{2}$. Indeed, we have
@@
Combining Steps 1-5, we obtain:
$\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}+\||y|v_{\text{re}}\|_{L^{2}}^{2}\leq C$
@@
3.2. Weighted $H^{2}$ Estimates for $v_{\textnormal{re}}$
With the weighted $H^{1}$ bounds established in Proposition 3.1, we now seek higher-order regularity for the remainder $v_{\text{re}}$. In the two-dimensional setting, $H^{1}$ regularity is insufficient to ensure a pointwise estimate. And it is necessary to establish estimates in the weighted Sobolev space $H^{2}$. The main goal of this section is to prove the $H^{2}$-estimate for $\sqrt{1+|y|^{2}}v_{\text{re}}$, which implies that $\sqrt{1+|y|^{2}}v_{\text{re}}(y)$ is bounded. We begin with a lemma establishing the improved regularity of a general weak solution in the usual Sobolev space.
@@ 3.2. Weighted $H^{2}$ Estimates for $v_{\textnormal{re}}$
$\begin{cases}-\Delta V+V\cdot\nabla V-\dfrac{1}{2}(y\cdot\nabla V+V)+\nabla P=f&\text{in }\mathbb{R}^{2},\\ \operatorname{div}V=0.\end{cases}$
@@
Then, there exists a universal positive constant $C$ such that
@@
First, note that
@@
$\|v_{\text{re}}\|_{H^{2}(\mathbb{R}^{2})}\leq C(A,\beta),$
and
$\|v_{\text{re}}\|_{L^{\infty}(\mathbb{R}^{2})}\leq C(A,\beta).$
@@
$f=-(v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{0})\in L^{2}(\mathbb{R}^{2}),~\|f\|_{\mathbb{R}^{2}}\leq C(A,\beta)$
@@
By the Sobolev embedding, we have
$\|v_{\text{re}}\|_{L^{\infty}(\mathbb{R}^{2})}\leq C(A,\beta).$
@@
Having established the $H^{1}$-estimate of $\sqrt{1+|y|^{2}}v_{\text{re}}$, we now focus on the second-order derivative estimates. First, we observe the following expansion:
@@
To ensure a rigorous derivation, one should choose $\varphi(y)=-D_{k}^{-h}(h_{\epsilon}^{2}D_{k}^{h}v_{\text{re}})=-D_{k}^{-h}h_{\epsilon}^{2}D_{k}^{h}v_{\text{re}}$ as a test function in (3.1), where $h_{\epsilon}(y)=\frac{|y|}{(1+\epsilon|y|^{2})^{1/2}}$, and subsequently pass to the limit $\epsilon\to 0$. For the sake of clarity in the following presentation, we perform the estimates using the formal weight $h(y)=|y|$ and the test function $\varphi(y)=-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}})$.
Below, we use $C$ to denote a constant depending on $A$ and $\beta$, which may vary from line to line. Testing (3.1) with $\varphi(y)=-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}})$, we obtain
@@
$-2\int_{\mathbb{R}^{2}}|y|D_{k}^{h}\nabla v_{\text{re}}:\left(\frac{y}{|y|}\otimes D_{k}^{h}v_{\text{re}}\right)dy$
$\leq$ $C\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$
@@
For the pressure term, a straightforward calculation yields:
@@
$=2\int_{\mathbb{R}^{2}}D_{k}^{h}qy\cdot(D_{k}^{h}v_{\text{re}})\,dy.$
$\leq C\|\nabla q\|_{L^{2}}^{2}+\frac{1}{20}\|yD_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}.$
@@
Similarly, for the interaction terms involving $v_{0}$, we have:
@@
$\leq$ $C\||y|^{2}\nabla^{2}v_{0}\|_{L^{\infty}}\|v_{\text{re}}\|_{L^{2}}\|\nabla v_{\text{re}}\|_{L^{2}}+C\||y|^{2}\nabla v_{0}\|_{L^{\infty}}\|\nabla v_{\text{re}}\|_{L^{2}}^{2}\leq C$
@@
3.3. Weighted $H^{2}$-Estimates for $\nabla v_{\textnormal{re}}$

### v2

@@ 2.3. Proof of the energy estimates
since we have from (2.24) that
${\rm div\thinspace}F(v_{0},w)={\rm div\thinspace}\left(-\Delta v_{1}-\frac{1}{2}v_{1}-\frac{1}{2}y\cdot\nabla v_{1}\right)=0,$
where ${\rm div\thinspace}v_{1}=0$ has been used. It follows from the Calderon–Zygmund estimates ([39, Chapter V]) that there exists a constant $c(p)$ such that
@@
$\qquad\qquad+\left.\|v_{1}\nabla v_{2}\|_{L^{p}}+\|v_{2}\nabla v_{2}\|_{L^{p}}\right).$
@@
We also have from (2.29) that
@@
Indeed, it follows from Lemma 2.2 that for any $1<p<2$, one has
$\int_{\mathbb{R}^{2}}|\Delta w|^{p}\\ \leq C(A,\beta,p),$
and that
Hence, using (2.24), we have for any $1<p<2$ that
$\int_{\mathbb{R}^{2}}|F(v_{0},w)|^{p}\leq$ $C\int_{\mathbb{R}^{2}}\left|2\nabla\eta_{R_{0}}\nabla v_{0}+\Delta\eta_{R_{0}}v_{0}+\frac{1}{2}y\cdot\nabla\eta_{R_{0}}v_{0}+\frac{1}{2}w+\frac{1}{2}y\cdot\nabla w\right|^{p}+\int_{\mathbb{R}^{2}}|\Delta w|^{p}$
$\leq$ $C(A,\beta,p),$
Now we can conclude by using (2.38) that
@@
The application of the Gagliardo–Nirenberg interpolation inequality for any fixed $1<p<2$ also gives that
@@
Throughout the proof, the finiteness of the Dirichlet energy of $v_{1}$ (or equivalently Dirichlet energy of $v_{0}$) (2.30) is the only place where we require the Hölder continuity exponent $\beta>0$.
We indeed have proved $H^{1}$-estimate (2.11) and $L^{p}$-estimate of $v_{\textrm{re}}$ when $1<p\leq 2$. To prove the rest of Theorem 2.1, we use the $H^{2}$-estimate of $v_{\text{re}}$ to be proved in Proposition 3.2. The proof of Proposition 3.2 only needs $H^{1}$-estimate (2.11) of $v_{\textrm{re}}$.
@@
Using the standard bootstrapping argument, along with the regularity estimates of the Stokes equations, one can show that $v_{\text{re}}\in H^{1}(\mathbb{R}^{2})$ is indeed smooth. For example, one can refer to [48, Theorem 3.8] for a detailed presentation of the bootstrapping argument. This also implies the smoothness of $u(x,1)$ as $e^{\Delta}u_{0}$ is already smooth. We next prove the corresponding estimates for $v_{\text{re}}$. Based on the definitions of $v_{\textrm{re}}$ and $v_{2}$ in (2.6) and (2.22) respectively, we have
$v_{\textrm{re}}-v_{2}=v_{1}-v_{0}=(\eta_{R_{0}}-1)v_{0}+w.$
It then follows from Lemmas 2.1 and 2.2 and Theorem 2.2 that for $1<p<2$,
@@
To show higher integrability, we show that the $H^{1}$-estimates for $v_{\textrm{re}}$ can be improved to $H^{2}(\mathbb{R}^{2})$, which will be done in Proposition 3.2 below. It follows that
@@
the Calderon–Zygmund estimates and the estimates for $v_{0}$ and $v_{\text{re}}$ give that for any $p>1$
@@
Similarly, the Calderon–Zygmund estimates (2.35), together with the above estimates of $v_{0}$ and $v_{\text{re}}$, yield that for any $p>1$
$\|\nabla q\|_{L^{p}(\mathbb{R}^{2})}\leq C(A,\beta,p).$
This finishes the proof of Theorem 2.1. ∎
@@ 3. Weighted Energy Estimates and Pointwise Decay
In this section, we establish the weighted energy estimate and the pointwise decay estimates for the remainder part $v_{\text{re}}$. For the three-dimensional case, Jia and Šverák [25] first utilized localized smoothing estimates for local Leray solutions to derive pointwise decay. Here, we adopt the methods developed in [33]. We first give the weighted energy estimates that $\sqrt{1+|y|^{2}}v_{\text{re}}$ in $H^{2}(\mathbb{R}^{2})$. This gives a first pointwise decay via Sobolev embeddings, i.e.,
@@
Furthermore, we prove the optimal decay for $v_{\text{re}}$ if the initial data $u_{0}\in C^{2}_{loc}$.
@@
(3.1) $\left\{\begin{aligned} &-\Delta v_{\text{re}}-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{\text{re}}\cdot\nabla v_{\text{re}}+\nabla q=-v_{0}\cdot\nabla v_{0},\\ &{\rm div\thinspace}v_{\text{re}}=0,\end{aligned}\right.\textrm{ on }\mathbb{R}^{2}.$
@@
3.1. Weighted $H^{1}$-Estimates of $v_{\text{re}}$
As a first step towards the higher-order estimates, we establish a weighted $H^{1}$ bound for the remainder term $v_{\text{re}}$. It can be seen in the proof that the convection term $\frac{1}{2}y\cdot\nabla v_{\text{re}}$ plays an important role.
@@ 3.1. Weighted $H^{1}$-Estimates of $v_{\text{re}}$
Under the same assumption of Theorem 2.1, consider the corresponding solution $(v_{\text{re}},q)\in X\times L^{2}(\mathbb{R}^{2})$ to (3.1), then we have
@@
$|\nabla(\sqrt{1+|y|^{2}}v_{\text{re}})|\leq\frac{|y|}{\sqrt{1+|y|^{2}}}|v_{\text{re}}|+\sqrt{1+|y|^{2}}|\nabla v_{\text{re}}|.$
Hence it holds that
@@
As we already have the $H^{1}$-estimates of $v_{\text{re}}$ due to Theorem 2.1, it suffices to prove the $H^{1}$-estimates of $|y|v_{\text{re}}$. The key is to test (3.1) with $\phi=|y|^{2}v_{\text{re}}$.
For a more rigorous derivation, one can use a regularized weight $h_{\epsilon}(y)=\frac{|y|}{(1+\epsilon|y|^{2})^{1/2}}$ and test (3.1) with $\phi=|h_{\epsilon}|^{2}v_{\text{re}}$, then pass to the limit $\epsilon\to 0$. For the sake of clarity, we present the estimate using the weight $h(y)=|y|$ and the proof using the weight $h_{\epsilon}(y)$ follows analogously with minor modification.
@@
$=\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}-2\|v_{\text{re}}\|_{L^{2}}^{2}$
$\geq\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}-C,$
where we have used (2.11).
Step 2: The Drift and Linear Damping Terms. The linear drift and damping part of the operator is $\mathcal{L}(v_{\text{re}})=-\frac{1}{2}v_{\text{re}}-\frac{1}{2}y\cdot\nabla v_{\text{re}}$. Testing this with $|y|^{2}v_{\text{re}}$, we compute that
@@
In 2D, we calculate the divergence explicitly to get
@@
Combining with the damping term $\frac{1}{2}v_{\text{re}}$ yields
@@
This positivity is specific to the special structure of the Leray equations and ensures the weighted $L^{2}$ norm is controlled.
@@
By Lemma 2.1, $|v_{0}(y)|\leq C(1+|y|)^{-1}$. Hence, together with (2.11), we have
@@
Step 4: Nonlinear term $v_{\text{re}}\cdot\nabla v_{\text{re}}$ and pressure term. For the nonlinear term $v_{\text{re}}\cdot\nabla v_{\text{re}}$, we have similarly as above that
@@
$\int_{\mathbb{R}^{2}}\nabla q\cdot|y|^{2}v_{\text{re}}$ $=-\int_{\mathbb{R}^{2}}2qv_{\text{re}}\cdot y\leq\frac{1}{10}\int_{\mathbb{R}^{2}}|y|^{2}|v_{\text{re}}|^{2}+C.$
Step 5: The source terms. Using Lemma 2.1 and (2.11), we have
@@
Combining Steps 1-5 gives
$\|\nabla(|y|v_{\text{re}})\|_{L^{2}}^{2}+\||y|v_{\text{re}}\|_{L^{2}}^{2}\leq\||y|\nabla v_{\text{re}}\|_{L^{2}}^{2}+\||y|v_{\text{re}}\|_{L^{2}}^{2}+\|v_{\text{re}}\|_{L^{2}}^{2}\leq C.$
@@
3.2. Weighted $H^{2}$-Estimates of $v_{\textnormal{re}}$
With the weighted $H^{1}$ bounds established in Proposition 3.1, we now seek higher-order regularity for the remainder $v_{\text{re}}$. In the two-dimensional setting, $H^{1}$ regularity is insufficient to ensure a pointwise estimate. And it is necessary to establish estimates in the weighted Sobolev space $H^{2}$. The main goal of this section is to prove the $H^{2}$-estimate for $\sqrt{1+|y|^{2}}v_{\text{re}}$, which implies that $\sqrt{1+|y|^{2}}v_{\text{re}}(y)$ is bounded. We begin with a lemma on the improved regularity of a general weak solution in the usual Sobolev space.
@@ 3.2. Weighted $H^{2}$-Estimates of $v_{\textnormal{re}}$
$\left\{\begin{aligned} &-\Delta V+V\cdot\nabla V-\dfrac{1}{2}(y\cdot\nabla V+V)+\nabla P=f,\\ &\operatorname{div}V=0,\end{aligned}\right.\textrm{ in }\mathbb{R}^{2}.$
@@
Then there exists a universal positive constant $C$ such that
@@
First, we note that
@@
$\|v_{\text{re}}\|_{H^{2}(\mathbb{R}^{2})}\leq C(A,\beta).$
@@
$f=-(v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{0})\in L^{2}(\mathbb{R}^{2}),~\textrm{and}~\|f\|_{L^{2}(\mathbb{R}^{2})}\leq C(A,\beta)$
@@
This finishes the proof of the proposition. ∎
@@
Having established the $H^{1}$-estimate of $\sqrt{1+|y|^{2}}v_{\text{re}}$, we now focus on the second-order derivative estimates. First, we observe the following expansion for $i,j=1,2$:
@@
For the sake of clarity for presentation, we perform the estimates using the weight $h(y)=|y|$ and the test function $\varphi(y)=-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}})$. The more rigorous proof employing $h_{\epsilon}(y)=\frac{|y|}{(1+\epsilon|y|^{2})^{1/2}}$ and passing $\epsilon\to 0$ follows analogously with minor modifications.
Below, we use $C$ to denote a constant depending on $A$ and $\beta$, which may vary from line to line. Testing (3.1) with $\varphi(y)=-D_{k}^{-h}(|y|^{2}D_{k}^{h}v_{\text{re}})$ ($k=1,2$), we obtain
@@
$-2\int_{\mathbb{R}^{2}}|y|D_{k}^{h}\nabla v_{\text{re}}:\left(\frac{y}{|y|}\otimes D_{k}^{h}v_{\text{re}}\right)dy\leq$ $C\|\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}\||y|D_{k}^{h}\nabla v_{\text{re}}\|_{L^{2}(\mathbb{R}^{2})}$
@@
For the pressure term, a straightforward calculation yields
@@
$=2\int_{\mathbb{R}^{2}}D_{k}^{h}qy\cdot(D_{k}^{h}v_{\text{re}})\,dy$
$\leq C\|\nabla q\|_{L^{2}}^{2}+\frac{1}{20}\|yD_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}$
$\leq C+\frac{1}{20}\|yD_{k}^{h}v_{\text{re}}\|_{L^{2}}^{2}.$
@@
Similarly, for the interaction terms involving $v_{0}$, we have
@@
$\leq$ $C\||y|^{2}\nabla^{2}v_{0}\|_{L^{\infty}}\|v_{\text{re}}\|_{L^{2}}\|\nabla v_{\text{re}}\|_{L^{2}}+C\||y|^{2}\nabla v_{0}\|_{L^{\infty}}\|\nabla v_{\text{re}}\|_{L^{2}}^{2}\leq C.$
@@
3.3. Weighted $H^{2}$-Estimates of $\nabla v_{\textnormal{re}}$